Electromagnetic radiation travels at a speed of 3.0 x 108 m/s.

a. What is the wavelength of a radio wave that has a frequency of 3.5 x105 Hz?

b. What is the frequency of an X-ray that has a wavelength of 2.5 x 10-10 m?


1
Expert's answer
2021-03-29T08:58:56-0400

"a.\ \ \ \lambda=c/\nu=3\cdot10^8/3.5\cdot10^5=857.1\ (m)"


"b. \ \ \nu=c/\lambda=3\cdot 10^8/2.5\cdot 10^{-10}=1.2\cdot 10^{18}\ \ (Hz)"